# Dichotomy theorem distinguishing non-integrability and the lowest-order Yang-Baxter equation for isotropic spin chains

### v1

@@ I.1 Yang-Baxter equation and the quantum inverse scattering method
$R_{12}(\lambda)R_{23}(\lambda+\mu)R_{23}(\mu)=R_{23}(\mu)R_{13}(\lambda+\mu)R_{12}(\lambda).$ (S.1)
@@
Plugging $I=R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{ab}(\lambda+\lambda^{\prime})$ to the left end of the product and applying the Yang-Baxter equation (S.1) repeatedly, we have
@@
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{ab}(\lambda+\lambda^{\prime})R_{1a}(\lambda)R_{1b}(\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{ab}(\lambda+\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)R_{ab}(\lambda+\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
@@
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda+\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)\cdots R_{Nb}(\lambda^{\prime})R_{Na}(\lambda)R_{ab}(\lambda+\lambda^{\prime})]$
@@ I.3 Boost operator
We first suppose that a Hamiltonian satisfies the Yang-Baxter equation. Then, the corresponding transfer matrix $T(\lambda)$ satisfies
$[B,T(\lambda)]=\frac{d}{d\lambda}T(\lambda)+O(\lambda^{2}),$ (S.20)
which is shown as follows. Let ${\cal S}=T(0)$ be the shift operator by one site. We expand $T(\lambda+\varepsilon)$ up to both $\lambda$ and $\varepsilon$ at the first order as
$T(\lambda+\varepsilon)={\cal S}\left(1+H(\lambda+\varepsilon)+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda\varepsilon+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\varepsilon+O(\lambda^{2},\varepsilon^{2})\right),$ (S.21)
which implies
$\frac{d}{d\lambda}T(\lambda)={\cal S}\left(H+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\right)+O(\lambda^{2}).$ (S.22)
On the other hand, $[B,T(\lambda)]$ up to the first order of $\lambda$ is calculated as
$[B,T(\lambda)]=$ $\sum_{j=1}^{N}[jh_{j,j+1},{\cal S}]+\sum_{i,j=1}^{N}[jh_{j,j+1},{\cal S}h_{i,i+1}]\lambda+O(\lambda^{2})$
$=$ ${\cal S}\left(H+\sum_{j=1}^{N}h_{j,j+1}^{2}\lambda+2\sum_{i<j}h_{i,i+1}h_{j,j+1}\lambda\right)+O(\lambda^{2}),$ (S.23)
which confirms Eq. (S.20).
@@
$[B,T^{m}(\lambda)]=\sum_{l=0}^{m}T^{l}(\lambda)[B,T(\lambda)]T^{l-m-1}(\lambda)=\sum_{l=0}^{m}T^{l}(\lambda)\frac{d}{d\lambda}T(\lambda)T^{l-m-1}(\lambda)+O(\lambda^{2})=\frac{d}{d\lambda}T^{m}(\lambda)+O(\lambda^{2}),$ (S.24)
@@
$[B,f(T(\lambda))]=\frac{d}{d\lambda}f(T(\lambda))+O(\lambda^{2})$ (S.25)
@@
This relation guarantees that $Q^{m}$ is indeed an $m$-local conserved quantity.
We next consider a general Hamiltonian $H$ on which the existence of the Yang-Baxter equation is not assumed. In this case, the obtained $m$-support operator $Q^{m}$ by Eq. (S.26) is no longer proven to be a conserved quantity.
@@
On the other hand, whether $Q^{m}$ with $m\geq 5$ are conserved quantities or not has not yet been proven. Grabowski and Mathieu conjectured as conjecture 2 in Ref. s-GM95-2 that if Eq. (S.28) holds, then this boost operator indeed gives an infinite sequence of local conserved quantities.

### v2

@@ I.1 Yang-Baxter equation and the quantum inverse scattering method
$R_{12}(\lambda)R_{13}(\lambda+\mu)R_{23}(\mu)=R_{23}(\mu)R_{13}(\lambda+\mu)R_{12}(\lambda).$ (S.1)
@@
Plugging $I=R^{-1}_{ab}(\lambda-\lambda^{\prime})R_{ab}(\lambda-\lambda^{\prime})$ to the left end of the product and applying the Yang-Baxter equation (S.1) repeatedly, we have
@@
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda-\lambda^{\prime})R_{ab}(\lambda-\lambda^{\prime})R_{1a}(\lambda)R_{1b}(\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda-\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{ab}(\lambda-\lambda^{\prime})R_{2a}(\lambda)R_{2b}(\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda-\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)R_{ab}(\lambda-\lambda^{\prime})\cdots R_{Na}(\lambda)R_{Nb}(\lambda^{\prime})]$
@@
$=$ $\mathrm{Tr}_{a,b}[R^{-1}_{ab}(\lambda-\lambda^{\prime})R_{1b}(\lambda^{\prime})R_{1a}(\lambda)R_{2b}(\lambda^{\prime})R_{2a}(\lambda)\cdots R_{Nb}(\lambda^{\prime})R_{Na}(\lambda)R_{ab}(\lambda-\lambda^{\prime})]$
@@ I.2 Reshetikin condition
The Reshetikin condition is thus the lowest-order nontrivial relation of the Yang-Baxter equation. This relation is sometimes employed to seek numerically new integrable models s-LMS25 .
@@ I.3 Boost operator
We first suppose that a Hamiltonian satisfies the Yang-Baxter equation (S.1). Then, the corresponding transfer matrix $T(\lambda)$ satisfies
$[B,T(\lambda)]=\frac{d}{d\lambda}T(\lambda),$ (S.20)
which is shown as follows s-Tel82 .
Differentiating the Yang-Baxter equation (S.1) by $\mu$ and setting $\mu=0$ with multiplying $\Pi_{23}$ from right, we find
$\left(\frac{d}{d\lambda}R_{12}(\lambda)\right)R_{13}(\lambda)-R_{12}(\lambda)\left(\frac{d}{d\lambda}R_{13}(\lambda)\right)=[R_{12}(\lambda)R_{13}(\lambda),h_{2,3}],$ (S.21)
where we used $R(0)=\Pi$ and $\left.\frac{d}{d\mu}R(\mu)\right|_{\mu=0}=\Pi h$. This relation is sometimes called the Sutherland equation. We reassign subscripts as $1\to a$, $2\to i$, $13\to i+$ 1, and multiply $\prod_{j\leq i-1}R_{ja}(\lambda)$ from left and $\prod_{j\geq i+2}R_{ja}(\lambda)$ from right, which leads to
$\prod_{j\leq i-1}R_{ja}(\lambda)\left(\frac{d}{d\lambda}R_{ia}(\lambda)\right)\prod_{j\geq i+1}R_{ja}(\lambda)-\prod_{j\leq i}R_{ja}(\lambda)\left(\frac{d}{d\lambda}R_{i+1,a}(\lambda)\right)\prod_{j\geq i+2}R_{ja}(\lambda)=[\prod_{j}R_{ja}(\lambda),h_{i,i+1}].$ (S.22)
Finally, we multiply $-i$ and take the sum with respect to $i$, which reads
$\frac{d}{d\lambda}\prod_{j}R_{ja}(\lambda)=[B,\prod_{j}R_{ja}(\lambda)].$ (S.23)
Here we used the telescope-type trick. Ignoring the boundaries and taking its trace with $a$, we arrive at Eq. (S.20).
@@
$[B,T^{m}(\lambda)]=\sum_{l=0}^{m}T^{l}(\lambda)[B,T(\lambda)]T^{l-m-1}(\lambda)=\sum_{l=0}^{m}T^{l}(\lambda)\frac{d}{d\lambda}T(\lambda)T^{l-m-1}(\lambda)=\frac{d}{d\lambda}T^{m}(\lambda),$ (S.24)
@@
$[B,f(T(\lambda))]=\frac{d}{d\lambda}f(T(\lambda))$ (S.25)
@@
This relation guarantees that $Q^{m}$ is indeed an $m$-local quantity.
We next consider a general Hamiltonian $H$ on which the existence of the Yang-Baxter equation is not assumed. In this case, the obtained $m$-support operator $Q^{m}$ by Eq. (S.26) is no longer proven to be a conserved quantity at present.
@@
The conservation of $Q^{m}$ with $m\geq 5$ has recently been solved by Hokkyo:
Lemma S.1 (Hokkyo s-Hok25-2 ).
Consider a shift-invariant nearest-neighbor interaction Hamiltonian $H=\sum_{i}h_{i,i+1}$. Suppose that $Q^{3}=[B,H]$ is a conserved quantity; $[Q^{3},H]=0$. Then, $Q^{m}$ given by Eq. (S.26) are also conserved quantities; $[Q^{m},H]=0$.
We here draw the outline of the proof of Lemma. S.1. We prove $[Q^{m},H]=0$ by induction on $m$. We first show that $Q^{m}$ is shift-invariant. Then, using the Jacobi identity we have
$[H,[H,Q^{m+1}]]=[H,[H,[B,Q^{m}]]]=[H,[Q^{m},[B,H]]]=-[Q^{m},[Q^{3},H]]-[Q^{3},[H,Q^{m}]]=0.$ (S.30)
Finally, we notice that $[H,[H,Q^{m+1}]]=0$ implies $[H,Q^{m+1}]=0$, since the action of $[H,\cdot\ ]$ maps nonzero off-diagonal elements with respect to the eigenstates of $H$ onto nonzero ones (i.e., if $Q^{m+1}$ has a nonzero off-diagonal element, then the same off-diagonal element of $[H,Q^{m+1}]$ is still nonzero, implying $[H,[H,Q^{m+1}]]\neq 0$). This completes the proof.
(We note that Zhang s-Zha25 also claims the same statement as Lemma. S.1. However, as the authors have examined, this proof has a logical gap.)